# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

### v1

@@ 7.2 First characterization
We start with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
@@
Lemma 7.3 (Conical projections yield facial subspaces).
@@
Since $K$ is finite dimensional, there is a finite number of vectors $v_{1},\ldots,v_{m}\in K$ such that $\operatorname{\mathrm{span}}(K)=\operatorname{\mathrm{span}}(v_{1},\ldots,v_{m})$. Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Lemma 6.3, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a facial subspace, which is the same set as $U\cap K^{\perp}$. ∎
Lemma 7.4 (Normal to the generated facial subspace of one convex set).
@@
Let $U=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. we first show that $U\in\mathcal{F}(C\cap D;C)$. Note that $T\in\mathcal{F}(C\cap D;C)$ by Lemma 6.10. Then we can use Lemma 7.3 because $N_{C}^{T}(C\cap T_{a})\subseteq N_{C}^{T}(C\cap D)$, which shows that $U\in\mathcal{F}(C\cap D;C)$ as well.
In the remainder of the proof we show that the assumption of Lemma 6.9 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
@@
Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Lemma 6.9. ∎
Proof of Proposition 4.1.
@@
Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Lemma 6.10, by definition of the generated facial subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
@@
From Proposition 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a facial subspace of $C$ that contains $C\cap D$, and so by Definition 6.5, $H_{C}(C\cap D)\subseteq T$. Therefore, for $^{\prime}:=H_{C}(C\cap D)+H_{D}(C\cap D), TT$$^{\prime}\subseteq T$.
For the other containment, we argue by contradiction: let $T^{\prime}\subsetneq T$. Then there is a unit vector $v\in T$ with $v\perp T^{\prime}$. Then $v\perp H_{C}(C\cap D)$, which by Lemma 7.4, implies that $v\in N_{C}^{T}(C\cap T)$. By Lemma 6.2, $v\in N_{C}^{T}(C\cap T)$ as well. And since $N_{C}^{T}(C\cap T)\subseteq N_{C}^{T}(C\cap D)$, we find that $v,-v\in N_{C}^{T}(C\cap D)$. By a similar argument, $v,-v\in N_{D}^{T}(C\cap D)$. But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Lemma 6.10 means that $v\in N_{C-D}^{T}(0)$. But by Lemma 6.9, $T$ cannot contain such a nested normal because by Definition 6.6, $T=H_{C-D}(0)$. Therefore, by contradiction $T=T^{\prime}$. ∎
@@ 7.3 Second characterization
First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Lemma 6.10. Applying Lemma 7.3 at each step, we find that $\forall i\in\{0,\ldots,\ell\},\,T_{i}\in\mathcal{F}(C-D;S)$. Finally, note that whenever $N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)]\neq\{0\}$, $\dim(T_{i+1})\leq\dim(T_{i})-1$. Therefore, the dimension may be reduced at most $n$ times, which implies that $\ell\leq n$. On the other hand, the fact that $T_{\ell}=T_{\ell+1}$ implies that $N^{T_{\ell}}_{C-D}(S)=\{0\}$. Then from Lemma 6.9 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.6. ∎
7.4 Aknowledgements
The author would like to thank Sharvaj Kubal, prof. Yaniv Plan, and prof. Ozgur Yilmaz for many fruitful discussions. A special thanks to prof. Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual would be a promising avenue of research. This paper was produced with the Latex Exporter plugin in Obsidian.
range
pages36 rangepages12 rangepages13 rangepages12 rangepages25 rangepages17 rangepages34

### v3

@@ 7.2 Analytical characterizations
We begin with the containment ($\supseteq$). Let $x\in N_{C}^{V}(S).$ Then $\forall s\in S,y\in C\cap(V+S)$, note that $y-s\in U$, and so
@@
Lemma 7.3 (Conical projections yield supporting subspaces).
@@
Since $K$ is finite dimensional, there is a finite number of vectors $v_{1},\ldots,v_{m}\in K$ such that $\operatorname{\mathrm{span}}(K)=\operatorname{\mathrm{span}}(v_{1},\ldots,v_{m})$. Then since $v_{1},\ldots,v_{m}\in N_{C}^{U}(S)$, by repeated application of Theorem 6.3, $U\cap v_{1}^{\perp}\ldots\cap v_{m}^{\perp}$ is a supporting subspace, which is the same set as $U\cap K^{\perp}$. ∎
Lemma 7.4 (Generated supporting subspace of one of the two convex sets).
@@
Let $U=T\cap N_{C}^{T}(C\cap T_{a})^{\perp}$. we first show that $U\in\mathcal{F}(C\cap D;C)$. Note that $T\in\mathcal{F}(C\cap D;C)$ by Theorem 6.6. Then we can use Lemma 7.3 because $N_{C}^{T}(C\cap T_{a})\subseteq N_{C}^{T}(C\cap D)$, which shows that $U\in\mathcal{F}(C\cap D;C)$ as well.
In the remainder of the proof we show that the assumption of Theorem 6.5 holds for $U$, namely that $N_{C}^{U}(C\cap D)=\{0\}$.
@@
Then $N_{C}^{U}(C\cap D)=\{0\}$, and so the result follows from Theorem 6.5. ∎
We are now in a position to prove the analytical characterizations in Section 4.
Proof of Lemma 4.1.
@@
Additionally, since we know that $T\in\mathcal{F}(C\cap D;C)$ from Theorem 6.6, by definition of the generated supporting subspace we know that $H_{C}(C\cap D)\subseteq T$. Therefore,
@@
From Lemma 4.1, we have that $T\in\mathcal{F}(C;C\cap D)$ is a supporting subspace of $C$ that contains $C\cap D$, and so by Definition 6.5, $H_{C}(C\cap D)\subseteq T$. Therefore, for $^{\prime}:=H_{C}(C\cap D)+H_{D}(C\cap D), TT$$^{\prime}\subseteq T$.
For the other containment, we argue by contradiction: let $T^{\prime}\subsetneq T$. Then there is a unit vector $v\in T$ with $v\perp T^{\prime}$. Then $v\perp H_{C}(C\cap D)$, which by Lemma 7.4, implies that $v\in N_{C}^{T}(C\cap T)$. By Lemma 6.2, $v\in N_{C}^{T}(C\cap T)$ as well. And since $N_{C}^{T}(C\cap T)\subseteq N_{C}^{T}(C\cap D)$, we find that $v,-v\in N_{C}^{T}(C\cap D)$. By a similar argument, $v,-v\in N_{D}^{T}(C\cap D)$. But then $v\in N_{C}^{T}(C\cap D)\cap(-N_{D}^{T}(C\cap D))$, which by Theorem 6.6 means that $v\in N_{C-D}^{T}(0)$. But by Theorem 6.5, $T$ cannot contain such a nested normal because by Definition 6.6, $T=H_{C-D}(0)$. Therefore, by contradiction $T=T^{\prime}$. ∎
Next, we present the proofs for the corollaries of Theorem 4.1, starting with a simple lemma.
Lemma 7.5 (Formulation of the tangent space).
For a convex set $C\subseteq\mathbb{R}^{n}$ and $S\subseteq C$ non-empty,
$\operatorname{\mathrm{span}}(C-C)=\operatorname{\mathrm{span}}(C-S).$
Proof of Lemma 7.5.
That $\operatorname{\mathrm{span}}(C-C)\supseteq\operatorname{\mathrm{span}}(C-S)$ is trivial, so we only show $\operatorname{\mathrm{span}}(C-C)\subseteq\operatorname{\mathrm{span}}(C-S)$. It suffices to show that $C-C\subseteq\operatorname{\mathrm{span}}(C-S)$; let $u,v\in C$, then for any fixed $x\in S$,
$u-v=(u-x)-(v-x)\in\operatorname{\mathrm{span}}(C-S).$
Proof of Corollary 4.3.
In this proof, we denote $F_{C}:=F_{C}(C\cap D)$ and $F_{D}:=F_{D}(C\cap D)$.
$\operatorname{\mathrm{span}}(F_{C}-F_{D})$ $=\operatorname{\mathrm{span}}(F_{C}-F_{D})+(C\cap D-C\cap D)$ (19)
$=\operatorname{\mathrm{span}}(F_{C}-F_{D}+(C\cap D-C\cap D))$ (20)
$=\operatorname{\mathrm{span}}((F_{C}-C\cap D)-(F_{D}-C\cap D))$ (21)
$=\operatorname{\mathrm{span}}(F_{C}-C\cap D)+\operatorname{\mathrm{span}}(F_{D}-C\cap D)$ (22)
$=\operatorname{\mathrm{span}}(F_{C}-F_{C})+\operatorname{\mathrm{span}}(F_{D}-F_{D}),$ (23)
$=H_{C}(C\cap D)+H_{D}(C\cap D).$ (24)
where Equation 19 holds because
$(C\cap D-C\cap D)\subseteq F_{C}-F_{D}\subseteq\operatorname{\mathrm{span}}(F_{C}-F_{D}),$
and Equation 23 follows from Lemma 7.5. The result then follows from Theorem 4.1. ∎
Proof of Corollary 4.2.
Throughout this proof, we omit the argument $C\cap D$, i.e., $H_{C}:=H_{C}(C\cap D)$, and similarly for $H_{D},F_{C},F_{D}$. We also define $\bar{F}_{C}=F_{C}-C\cap D$ and $\bar{F}_{D}=F_{D}-C\cap D$. By Lemma 7.5,
$H_{C}=\operatorname{\mathrm{span}}(F_{C}-F_{C})=\operatorname{\mathrm{span}}(\bar{F}_{C}).$
With this, starting with $T_{a}$ as in Theorem 4.1,
$T_{a}(C,D)$ $=H_{C}+H_{D}+C\cap D$
$=\operatorname{\mathrm{span}}(\bar{F}_{C})+\operatorname{\mathrm{span}}(\bar{F}_{D})+C\cap D$
$=\operatorname{\mathrm{span}}(\operatorname{\mathrm{span}}(\bar{F}_{C})\cup\operatorname{\mathrm{span}}(\bar{F}_{D}))+C\cap D$
$=\operatorname{\mathrm{span}}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D})+C\cap D$
$=\operatorname{aff}(\bar{F}_{C}\cup\bar{F}_{D}+C\cap D)$
$=\operatorname{aff}(F_{C}\cup F_{D}).$
@@ 7.3 Second characterization
First, $(N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)])=N^{T_{i}}_{C-D}(S)$ by Theorem 6.6. Applying Lemma 7.3 at each step, we find that $\forall i\in\{0,\ldots,\ell\},\,T_{i}\in\mathcal{F}(C-D;S)$. Finally, note that whenever $N^{T_{i}}_{C}(S)\cap[-N^{T_{i}}_{D}(S)]\neq\{0\}$, $\dim(T_{i+1})\leq\dim(T_{i})-1$. Therefore, the dimension may be reduced at most $n$ times, which implies that $\ell\leq n$. On the other hand, the fact that $T_{\ell}=T_{\ell+1}$ implies that $N^{T_{\ell}}_{C-D}(S)=\{0\}$. Then from Theorem 6.5 it follows that $T_{\ell}=H_{C-D}(S)$, which is $T(C,D)$ from Definition 6.6. ∎
7.4 Acknowledgements
The author would like to thank Sharvaj Kubal, Naomi Graham, prof. Yaniv Plan, and prof. Ozgur Yilmaz for many fruitful discussions. A special thanks to prof. Michael Friedlander for pointing out that a generalization of the Fenchel-Rockafellar dual and optimality conditions would be a promising research direction. This paper was produced with the Latex Exporter plugin in Obsidian.